True or false and give reason for your answer


i) If all the Eigen values of a matrix are real, then the
matrix must be symmetric.
ii) If V and W are real vector spaces, with T : V --> W
being a linear transformation then V/kerT isomorphic to W
iii) There are at least three different unitary matrices of order 2.
iv) There are two subspaces U and W of R3 such that U(intersection) W is empty.
v) If the determinant of a matrix is zero, then the matrix cannot be diagonalised.
1
Expert's answer
2020-12-25T15:21:52-0500

i) False

Reason:-

Let a matrix "A = \begin{bmatrix}1&2\\0&0\end{bmatrix}"

Clearly A is not symmetric matrix.

Now let "\lambda" be an eigen value of A.

So, "det(\lambda I-A)=0" "\Rightarrow \lambda=0" or, 1

"\therefore" Eigen values of A are real but A is not symmetric.


ii) False

Reason:-


Let "V=W=R^2"

And "T: V\to W" such that "T(\begin{bmatrix}a\\b\end{bmatrix} ^T)=\begin{bmatrix}a\\0\end{bmatrix} ^T"

It is easy to see that "T" is well-defined and "T" is a Linear Transformation.

[As "v_1=v_2 \Rightarrow T(v_1)=T(v_2)" and "T(v_1+cv_2)=T(v_1)+cT(v_2)"]


Now "Im(T) = \{T(v) : v\in V\} = \{\begin{bmatrix}a\\0\end{bmatrix}^T : a" is a real number"\}"

Clearly "Im(T) \subset W \Rightarrow |Im(T)|<|W|"


Now, From "1st" isomorphism theorem,

"V/ker(T) \cong Im(T)"

"\Rightarrow |V/ker(T)| = |Im(T)| < |W|"

"\Rightarrow |V/ker(T)| < |W|"


"\exists" no bijection between "V\/ker(T)" and "W"

So, "V\/ker(T)" is not isomorphic to "W"


iii) True

Reason :-

Let 3 matrices of order 2 are A, B and C, where

"A = \begin{bmatrix}1&0\\0&1\end{bmatrix}" , "B = \begin{bmatrix}0&i\\i&0\end{bmatrix}" and "C = \begin{bmatrix}i/2&\sqrt{3}/2\\\sqrt{3}/2&i/2\end{bmatrix}"

It is easy to check, "A(A^\theta)^T=B(B^\theta)^T=C(C^\theta)^T=I"

Where "M^\theta" is conjugate of M.

"\therefore A,B,C" are Unitary.


So, A,B,C are 3 different unitary matrices of order 2.

"\therefore \exists" at least 3 different unitary matrices of order 2.


iv) False

Reason :-

Let "U, W" are two subspaces of "\R^3" .

let "u\in U" and c be a scalar. "\Rightarrow cu\in U"

Take "c=0" "\therefore 0\in U"

Similarly "0\in W"

So, "0\in U\cap W"

"\therefore" "U\cap W" is not empty


v) False

Reason :-

Let "A = \begin{bmatrix}1&2\\0&0\end{bmatrix}" be a matrix of order 2.

Clearly "det(A)=0"


Let "P = \begin{bmatrix}0&1\\1&2\end{bmatrix}"


"\therefore P^{-1} = \begin{bmatrix}-2&1\\1&0\end{bmatrix}"


Now "PAP^{-1}= \begin{bmatrix}0&1\\1&2\end{bmatrix}\begin{bmatrix}1&2\\0&0\end{bmatrix}\begin{bmatrix}-2&1\\1&0\end{bmatrix}=\begin{bmatrix}0&0\\0&1\end{bmatrix}"

"\therefore PAP^{-1}" is a diagonal matrix. "\Rightarrow A" is diagonalizable.


So, the Determinant of "A" is but "A" is diagonalizable.